OpenStudy (anonymous):

1 - 2/3y = 9/5 - y/5 + 3/5 Solve for y

OpenStudy (anonymous):

hello and I hope your Sunday is well, what do you not understand I will help you.

OpenStudy (anonymous):

Thank you, I know you have to find the LCD, but I don't understand what to do after that

OpenStudy (anonymous):

okay yes you can find LCD, although there is one thing can be done before that. do you what can be done first?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

what? you tell me and I will say right or wrong..

OpenStudy (anonymous):

Oh, no I don't know what to do, I thought you meant do you want to know

OpenStudy (anonymous):

hint it has to do with Y..

OpenStudy (anonymous):

solve for y?

OpenStudy (anonymous):

close think a like terms..

OpenStudy (anonymous):

combine them?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

How can you combine 2/3y and y/5 though?

OpenStudy (anonymous):

never mind I thought the y was able be put togther on the five sorry.

OpenStudy (anonymous):

Oh, no problem, I have an easier problem I just don't understand one step I can draw it for you and could you possibly tell me why I'm doing this one step?

OpenStudy (anonymous):

so we have this math problem 1 - 2/3y = 9/5 - y/5 + 3/5 and we are going to solve for Y in because have factions we need to find an LCD for 9/5-3/5 what is that? I am taking this step by step with you

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now we have 1-2/3y=15-y/5

OpenStudy (anonymous):

why did the 3/5 turn into 15

OpenStudy (anonymous):

because you told me 15, right?

OpenStudy (anonymous):

the LCD is 15

OpenStudy (anonymous):

oh okay go one from that point.

OpenStudy (anonymous):

do you put 15 under the 1?

OpenStudy (anonymous):

1/15 yes

OpenStudy (anonymous):

ok so 1/15- 2y/15 = 9/15 - y/15 + 3/15?

OpenStudy (anonymous):

okay now we need to get y alone how?

OpenStudy (anonymous):

I am not sure

OpenStudy (anonymous):

look at -2y/15+ y/5 and the reason it is a plus because it crossed the equals so the sign goes from - to +

OpenStudy (anonymous):

it was y/15 forgot the one

OpenStudy (anonymous):

because we have -2y/15 +y/15, didive by y/15 so you will have -2y/15 /y/15 now do you see what can happen now?

OpenStudy (anonymous):

hey I will not give answers sorry..

OpenStudy (anonymous):

okay so I guess since are you off line I am no longer needed. Good luck